time period of an e- moving in 2nd bohr orbit is x.then the time period of the e- moving in 5th orbit of the same atom is ?

Grade:11

1 Answers

keshav agarwal
6 years ago
as atom is same in both cases so Z it would not change ‘Time Period’ of diffrent orbits
T=2(pie)r/v
which implies t is propotional to r/v
=> T is propotional to n^2/n^-1=n^3
=>T1/T2=n1^3/n2^3 = (2/5)^3= 8/25
=>T1/T2=8/25
please upvote my answer